OpenStudy (celinegirl):

Simplify 5 square root of 11 end root minus 12 square root of 11 end root minus 2 square root of 11. negative 33 square root of 9 negative 11 square root of 9 negative 9 square root of 11 negative 9 square root of 33

OpenStudy (johnweldon1993):

When we have the same square root across multiple terms...we just ignore the root and focus on the coefficients \[\large 5\sqrt{11} - 12\sqrt{11} - 2\sqrt{11}\] \[\large (5 - 12 - 2)\sqrt{11}\] \[\large -9\sqrt{11}\]

OpenStudy (celinegirl):

@johnweldon1993 thank you I was thinking it was C but i wanted to make sure (:

OpenStudy (johnweldon1993):

No problem! And since you had it right anyways...good job :D

OpenStudy (celinegirl):

@johnweldon1993 thanks (: